# MIT-CTP/5723

### v2

@@ 2 Background and related work
In Fig. 1 we sketch the typical data-analysis pipeline in particle physics. Detector measurements and theory simulations are processed in parallel and ultimately compared. Both use as a central representation the notion of particles. A particle is characterized by a discrete type label, an energy $E\in\mathbb{R}$, and a spatial momentum $\vec{p}\in\mathbb{R}^{3}$. Types include fundamental particles like electrons, photons, quarks, and gluons, composite particles like protons, as well as reconstructed objects like “jets” [69] or “particle-flow candidates” [70], which are the outputs of complex reconstruction algorithms. The energy and spatial momentum of a particle are conveniently combined into a four-momentum $p=(E,\vec{p})$.
@@
Conditional flow matching [52] is a simple and scalable training algorithm for continuous normalizing flows that does not require the simulation of trajectories during training. Instead, the objective is to match a vector field $v_{t}(x)$, parametrized by a neural network, onto a conditional target vector field $u_{t}(x|x_{1})$ along a conditional probability path $p_{t}(x|x_{1})$, minimizing the loss $\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],x_{1}\sim q(x_{1}),x\sim p_{t}(x|x_{1})}\lVert v_{t}(x)-u_{t}(x|x_{1})\rVert^{2}$, where $x_{1}\sim q(x_{1})$ are samples from the base distribution.
@@ 3.1 Lorentz-equivariant architecture
Any linear map $\phi:\mathbb{G}_{1,3}\to\mathbb{G}_{1,3}$ that is equivariant to $\mathrm{SO}^{+}(1,3)$ is of the form
$\phi(x)=\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}+\sum_{k=0}^{4}\,w_{k}e_{0123}\langle x\rangle_{k}$ (1)
@@
$\mathrm{Attention}(q,k,v)_{i^{\prime}c^{\prime}}=\sum_{i}\mathrm{Softmax}_{i}\Biggl{(}\frac{\sum_{c}\langle q_{i^{\prime}c},k_{ic}\rangle}{\sqrt{16n_{c}}}\Biggr{)}\;v_{ic^{\prime}}\,,$ (2)
@@
$\mathrm{LayerNorm}(x)=x\Big{/}{\sqrt{\frac{1}{n}\sum_{c=1}^{n}\sum_{k=0}^{4}\biggl{\lvert}\Bigl{\langle}\langle x_{c}\rangle_{k},\langle x_{c}\rangle_{k}\Bigr{\rangle}\biggr{\rvert}+\epsilon}}\,,$ (3)
@@ 3.2 Lorentz-equivariant flow matching
Concretely, reconstructed particles $p=(E,\vec{p})$ are often required to satisfy constraints of the form $p_{1}^{2}+p_{2}^{2}\geq p_{T\,\text{min}}^{2}$ and $p^{2}>0$. Following Refs. [16, 17, 44, 47], we parameterize this manifold with physically motivated coordinates $y=(y_{m},y_{p},\eta,\phi)$. The map to four-momenta is given by
@@
where $m^{2}=\exp(y_{m})$ and $p_{T}=p_{T,\text{min}}+\exp(y_{p})$. We define a constant diagonal metric in the coordinates $y$ and use the corresponding geodesics as probability paths. This Riemannian manifold is geodesically convex, meaning any two points are connected by a unique geodesic, and geodesically complete, meaning that paths thus never enter four-momentum regions forbidden by the phase-space cuts. By also running the ordinary differential equation (ODE) solver in these coordinates, we guarantee that each sample satisfies the four-momentum constraints. As an added benefit, this choice of metric compresses the high-energy tails of typical particle distributions and thus simplifies learning them correctly.
@@ 4.2 Top tagging
Next, we turn to the problem of classifying whether a spray of reconstructed hadrons originated from the decay of a top quark or any other process. This problem of top tagging is an important filtering step in any analysis that targets the physics of top quarks, the heaviest elementary particle in the Standard Model. We use the established top tagging dataset by Kasieczka et al. [49, 48] as a benchmark and compare to the published results for many algorithms and architectures.
@@ Results
As shown in Tbl. 1, L-GATr achieves a performance that is for all practical purposes on par with the strongest baselines.
@@ 4.3 Generative modelling
We train continuous normalizing flows based on an L-GATr network with the Riemannian flow matching objective described in Sec. 3. As baselines, we consider similar flow matching models, but use MLP and Transformer networks as score models, as proposed by Refs. [17, 44].
@@
Next, we evaluate the generative models quantitatively through the log likelihood of data samples under the trained models; see Appendix C for details. The left panel of Fig. 5 shows that the L-GATr models outperform both baselines across all different jet multiplicities. They maintain this performance advantage also for smaller training data size, as shown in the right panel.
@@ 4.4 Computational cost and scalability
Figure 5: Generative modelling: negative log likelihood on the test set (lower is better). Left: For different processes. Right: As a function of the training dataset size. We show the mean and standard deviation of three random seeds. The L-GATr flow outperforms the baselines in all processes and all training set sizes.
@@ Acknowledgements
We would like to thank Taco Cohen and Anja Butter for fruitful discussions, Nathan Hütsch for help with conditional flow matching, and David Ruhe for help with the CGENN code.
J.S., V.B. and T.P. are supported by the Baden-Württemberg-Stiftung through the program Internationale Spitzenforschung, project Uncertainties — Teaching AI its Limits (BWST_IF2020-010), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery, and through Germany’s Excellence Strategy EXC 2181/1 – 390900948 (the Heidelberg STRUCTURES Excellence Cluster). J.S. is funded by the Carl-Zeiss-Stiftung through the project Model-Based AI: Physical Models and Deep Learning for Imaging and Cancer Treatment. V.B. is supported by the BMBF Junior Group Generative Precision Networks for Particle Physics (DLR 01IS22079). J.T. is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/), by the U.S. Department of Energy Office of High Energy Physics under grant number DE-SC0012567, and by the Simons Foundation through Investigator grant 929241.
@@ Appendix A Geometric algebra
The fact that $\rho$ is an algebra homomorphism is equivalent to saying that the geometric product is equivariant to $\mathrm{Pin}(V)$. Furthermore, the grades in a geometric algebra form subrepresentations [13, Prop. 2]. Thus, the grade projections are equivariant. The pseudoscalar is a one-dimensional real representation, and thus must be invariant to any connected subgroup of $\mathrm{Pin}(V)$. Therefore, multiplying by the pseudoscalar is equivariant to the connected group $\mathrm{SO}^{+}(1,3)$. Hence, the linear layer in Eq. (1) is equivariant, for any value of the parameters $v,w$.
@@ Appendix B Architectures
We build a new architecture based on the Deep Sets framework [76] for tackling the amplitude surrogate task. Deep Sets is a permutation-invariant architecture that applies the same function to each element of an input set, aggregates the results with a permutation-invariant operation like a sum, and processes the outputs with another function.
Our adaptation for the amplitude regression tasks applies the Deep Sets approach to each subset of identical particles in the inputs, as the amplitudes are manifestly invariant under permutations of the four-momenta of particles of the same type. We thus apply a different preprocessing to each particle type and aggregate separately for each particle type. In addition to the particle-specific latent space samples, the input to our main network also includes the momentum invariants for all the particles involved in the process. DSI thus combines a Lorentz-equivariant, permutation-equivariant path with a non-Lorentz-equivariant non-permutation-equivariant path, allowing the network to learn whether to rely on equivariant or non-equivariant features.
@@
To ablate to what extent L-GATr’s performance is due to the geometric algebra representations and its equivariance properties and to what extent due to the Transformer architecture, we use L-GATr’s MLP block as a standalone network. We call this “Geometric Algebra Perceptron” (GAP). Instead of structuring our data as a set of tokens, we format the particle data as a single list of channels. This implies that interactions between particles within the model will be carried out in the linear layers, as opposed to the attention mechanism we use in L-GATr.
@@
Orthogonally to GAP, we also use a vanilla Transformer as a baseline in our experiments. We use a per-LayerNorm Transformer with multi-head attention and GELU nonlinearities. This setup mirrors that of L-GATr as closely as possible.
@@ C.1 Surrogates for QFT amplitudes
We generate training and evaluation data consisting of phase space inputs and their corresponding interaction amplitudes for processes $q\bar{q}\to Z+ng$, $n=\{1,4\}$, where an initial quark-antiquark pair interact to produce a $Z$ boson and a variable number of gluons. The amplitudes are invariant under the permutation of any identical particles and under Lorentz transformations. These datasets are generated by the MadGraph Monte Carlo event generator [3] in two steps. [^9] Available at https://launchpad.net/mg5amcnlo under a UoI-NCSA open source license. First, we use a standard run to generate the phase space distributions. This standard run applies importance sampling to produce unweighted samples, that is, events that are distributed according to the probability distribution that describes the physical interactions. Second, we re-compute the amplitude values corresponding to these phase-space samples with MadGraph’s standalone module.
@@
For the L-GATr model, we embed each particle as a token. It is characterized with its four-momentum, embedded as a grade-1 multivector in the geometric algebra, as well as a one-hot embedding of the particle type, embedded as a scalar. We standardize the four-momentum inputs to unit variance, using the same normalization for each component in order to not break Lorentz equivariance. In addition to the particle tokens we use one “global” token, initialized to zero. After processing these inputs with an L-GATr network, we select the scalar component of the global token and identify it as the amplitude output. We use 8 attention blocks, 32 multivector and 32 scalar channels, and 8 attention heads, resulting in $1.8\times 10^{6}$ learnable parameters.
For the CGENN, we minimally alter the graph neural network version of the model built for the top tagging classification task so that it is able to perform amplitude regression. We keep the hyperparameters proposed by [67] and use 72 hidden node features, 8 hidden edge features, and 4 blocks. This model features around $3.2\times 10^{5}$ trainable parameters. Scaling up the size of the CGENN so that it matches the parameter count in our L-GATr build does not translate into any visible improvement in the amplitude predictions.
@@
All models are trained by minimizing a mean squared error (MSE) loss on the preprocessed amplitude targets and by making use of the Adam optimizer. We use a batch size of 256 and a fixed learning rate of $10^{-4}$ for all baselines except CGENN, which we train with a batch size of 32 due to memory restrictions and the Cosine Annealing scheduler [55] with a maximum learning rate of $10^{-4}$. As for the number of training steps, MLP and DSI are trained for around $2.5\times 10^{6}$ iterations, the Transformer for around $10^{6}$ iterations and GAP, CGENN and GATr for $2.5\times 10^{5}$ iterations. We use no regularization method for any of our baselines except for CGENN, which includes by default a dropout of 0.2 at the output layer. We use early stopping across all training runs.
@@ C.2 Top tagging
As the top-tagging datasets operate with reconstructed particles as measured by a detector, we include the proton beam direction as an input to the network, which partially breaks the Lorentz equivariance of the process. We use additional channels with spacelike four-momentum $(0,0,0,1)$, but find that other choices lead to comparable results.
@@
L-GATr is trained by minimizing a binary cross entropy (BCE) loss on the top quark labels. We train it for $4\times 10^{5}$ steps using the EvoLved Sign Momentum (LION) [26] optimizer with a weight decay of 0.2 and a batch size of 128. We use a Cosine Annealing scheduler [55] with the maximum learning rate set at $3\times 10^{-4}$.

### v3

@@ 2 Background and related work
In Fig. 1 we sketch the typical data-analysis pipeline in particle physics. Its central idea is to take the data collected in the detectors as well as the predictions from different theories of physics, process both in parallel, and ultimately compare their predictions. The pipeline includes various steps, including the computation of scattering probabilities or amplitudes in mathematical frameworks called quantum field theories, the Monte-Carlo sampling from the theory, the simulated interaction of particles with the detector, the dimensionality reduction of the raw detector output to a small number of observables, the data filtering to extract only collisions of interest, and the statistical analysis of whether two predictions are consistent.
What most steps in this pipeline have in common is the notion of particles, the main representation of data in high-energy physics. A particle is characterized by a discrete type label, an energy $E\in\mathbb{R}$, and a spatial momentum $\vec{p}\in\mathbb{R}^{3}$. Types include fundamental particles like electrons, photons, quarks, and gluons, composite particles like protons, as well as reconstructed objects like “jets” [70] or “particle-flow candidates” [71], which are the outputs of complex reconstruction algorithms. The energy and spatial momentum of a particle are conveniently combined into a four-momentum $p=(E,\vec{p})$.
@@
Conditional flow matching [53, 3] is a simple and scalable training algorithm for continuous normalizing flows that does not require the simulation of trajectories during training. Instead, the objective is to match a vector field $v_{t}(x)$, parametrized by a neural network, onto a conditional target vector field $u_{t}(x|x_{1})$ along a conditional probability path $p_{t}(x|x_{1})$, minimizing the loss $\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],x_{1}\sim q(x_{1}),x\sim p_{t}(x|x_{1})}\lVert v_{t}(x)-u_{t}(x|x_{1})\rVert^{2}$, where $x_{1}\sim q(x_{1})$ are samples from the base distribution.
@@ 3.1 Lorentz-equivariant architecture
Any linear map $\mathrm{Linear}:\mathbb{G}_{1,3}\to\mathbb{G}_{1,3}$ that is equivariant to $\mathrm{SO}^{+}(1,3)$ is of the form
$\mathrm{Linear}(x)=\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}+\sum_{k=0}^{4}\,w_{k}e_{0123}\langle x\rangle_{k}$ (1)
@@
$\mathrm{Attention}(q,k,v)_{i^{\prime}c^{\prime}}=\sum_{i}\mathrm{Softmax}_{i}\Biggl{(}\sum_{c=1}^{n_{c}}\frac{\langle q_{i^{\prime}c},k_{ic}\rangle}{\sqrt{16n_{c}}}\Biggr{)}\;v_{ic^{\prime}}\,,$ (2)
@@
$\mathrm{LayerNorm}(x)=x\Big{/}{\sqrt{\frac{1}{n_{c}}\sum_{c=1}^{n_{c}}\sum_{k=0}^{4}\biggl{\lvert}\Bigl{\langle}\langle x_{c}\rangle_{k},\langle x_{c}\rangle_{k}\Bigr{\rangle}\biggr{\rvert}+\epsilon}}\,,$ (3)
@@ 3.2 Lorentz-equivariant flow matching
Concretely, reconstructed particles $p=(E,\vec{p})$ are often required to satisfy constraints of the form $p_{1}^{2}+p_{2}^{2}\geq p_{T\,\text{min}}^{2}$ and $p^{2}>0$. Following Refs. [17, 18, 45, 48], we parameterize this manifold with physically motivated coordinates $y=(y_{m},y_{p},\eta,\phi)$. These variables form an alternative basis for the particle four-momenta and are defined through the map
@@
where $m^{2}=\exp(y_{m})$ and $p_{T}=p_{T,\text{min}}+\exp(y_{p})$. This basis is better aligned with the physically relevant properties of particles in the context of a collider experiment: $\eta$ and $\phi$ represent the angle in which a particle is moving, $y_{p}$ is a measure of the momentum with which it moves away from the collision, and $y_{m}$ is related to its mass.
We define a constant diagonal metric in the coordinates $y$ and use the corresponding geodesics as probability paths. This Riemannian manifold is geodesically convex, meaning any two points are connected by a unique geodesic, and geodesically complete, meaning that paths thus never enter four-momentum regions forbidden by the phase-space cuts. By also running the ordinary differential equation (ODE) solver in these coordinates, we guarantee that each sample satisfies the four-momentum constraints. As an added benefit, this choice of metric compresses the high-energy tails of typical particle distributions and thus simplifies learning them correctly.
@@ 4 Experiments
We now demonstrate L-GATr in three applications. Each addresses a different problem in the data-analysis pipeline sketched in Fig. 1.
@@ 4.2 Top tagging
Next, we turn to the problem of classifying whether a spray of reconstructed hadrons originated from the decay of a top quark or any other process. This problem of top tagging is an important filtering step in any analysis that targets the physics of top quarks, the heaviest elementary particle in the Standard Model. Particle collisions involving these particles are of particular interest to physicists because the production and decay probabilities of top quarks are sensitive to several proposed theories of new physics, including for instance the existence of “supersymmetric” particles. We use the established top tagging dataset by Kasieczka et al. [50, 49] as a benchmark and compare to the published results for many algorithms and architectures.
@@ Results
As shown in Tbl. 1, L-GATr is on par with or better than even the strongest baselines on this well-studied benchmark.
@@ 4.3 Generative modelling
We train continuous normalizing flows based on an L-GATr network with the Riemannian flow matching objective described in Sec. 3. As baselines, we consider similar flow matching models, but use MLP and Transformer networks as score models, as proposed by Refs. [18, 45]. We also construct a flow matching model using the E(3)-equivariant GATr from Ref. [14]. Finally, we also train JetGPT [18] model, an autoregressive transformer architecture developed for particle physics that is not equivariant to the Lorentz symmetry.
@@
Next, we evaluate the generative models quantitatively through the log likelihood of data samples under the trained models; see Appendix C for details. The left panel of Fig. 5 shows that the L-GATr models outperform all baselines across all different jet multiplicities. They maintain this performance advantage also for smaller training data size, as shown in the right panel. The flow models, including L-GATr, are more data-efficient than the autoregressive transformer JetGPT.
@@ 4.4 Computational cost and scalability
Figure 5: Generative modelling: negative log likelihood on the test set (lower is better). Left: For different processes. Right: As a function of the training dataset size. We show the mean and standard deviation of three random seeds. The L-GATr conditional flow matching (CFM) model outperforms all other CFM models as well as the autoregressive transformer JetGPT, across all processes and all training set sizes.
@@ Acknowledgements
We would like to thank Taco Cohen and Anja Butter for fruitful discussions, Nathan Hütsch for help with conditional flow matching, and David Ruhe for support with the CGENN code.
J.S., V.B. and T.P. are supported by the Baden-Württemberg-Stiftung through the program Internationale Spitzenforschung, project Uncertainties — Teaching AI its Limits (BWST_IF2020-010), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery, and through Germany’s Excellence Strategy EXC 2181/1 – 390900948 (the Heidelberg STRUCTURES Excellence Cluster). J.S. is funded by the Carl-Zeiss-Stiftung through the project Model-Based AI: Physical Models and Deep Learning for Imaging and Cancer Treatment. V.B. is supported by the BMBF Junior Group Generative Precision Networks for Particle Physics (DLR 01IS22079). V.B. acknowledges financial support from the Grant No. ASFAE/2022/009 (Generalitat Valenciana and MCIN, NextGenerationEU PRTR-C17.I01). J.T. is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/), by the U.S. Department of Energy Office of High Energy Physics under grant number DE-SC0012567, and by the Simons Foundation through Investigator grant 929241.
@@ Appendix A Geometric algebra
The fact that $\rho$ is an algebra homomorphism is equivalent to saying that the geometric product is equivariant to $\mathrm{Pin}(V)$. Furthermore, the grades in a geometric algebra form subrepresentations [14, Prop. 2]. Thus, the grade projections are equivariant. The pseudoscalar is a one-dimensional real representation, and thus must be invariant to any connected subgroup of $\mathrm{Pin}(V)$. Therefore, multiplying by the pseudoscalar is equivariant to the connected group $\mathrm{SO}^{+}(1,3)$. Hence, the linear layer in Eq. (1) is equivariant, for any value of the parameters $v,w$.
@@ Appendix B Architectures
We build a new architecture based on the Deep Sets framework [77] that uses momentum invariants as part of the input for tackling the amplitude surrogate task. Deep Sets is a permutation-invariant architecture that applies the same function to each element of an input set, aggregates the results with a permutation-invariant operation like a sum, and processes the outputs with another function.
Our adaptation for the amplitude regression tasks applies the Deep Sets approach to each subset of identical particles in the inputs, as the amplitudes are manifestly invariant under permutations of the four-momenta of particles of the same type. We thus apply a different preprocessing to each particle type and aggregate them separately. In addition to the particle-specific latent space samples, the input to our main network also includes the momentum invariants for all the particles involved in the process. DSI thus combines a Lorentz-equivariant, permutation-equivariant path with a non-Lorentz-equivariant non-permutation-equivariant path, allowing the network to learn whether to rely on equivariant or non-equivariant features.
@@
To ablate to what extent L-GATr’s performance is due to the geometric algebra representations and its equivariance properties and to what extent due to the Transformer architecture, we use L-GATr’s MLP block as a standalone network. We call it “Geometric Algebra Perceptron” (GAP). Instead of structuring our data as a set of tokens, we format the particle data as a single list of channels. This implies that interactions between particles within the model will be carried out in the linear layers, as opposed to the attention mechanism we use in L-GATr.
@@
Orthogonally to GAP, we also use a vanilla Transformer as a baseline in our experiments. We use a pre-LayerNorm Transformer with multi-head attention and GELU nonlinearities. This setup mirrors that of L-GATr as closely as possible.
@@ C.1 Surrogates for QFT amplitudes
We generate training and evaluation data consisting of phase space inputs and their corresponding interaction amplitudes for processes $q\bar{q}\to Z+ng$, $n=\{1,2,3,4\}$, where an initial quark-antiquark pair interact to produce a $Z$ boson and a variable number of gluons. The amplitudes are invariant under the permutation of identical particles and under Lorentz transformations. These datasets are generated by the MadGraph Monte Carlo event generator [4] in two steps. [^9] Available at https://launchpad.net/mg5amcnlo under a UoI-NCSA open source license. First, we use a standard run to generate the phase space distributions. This standard run applies importance sampling to produce unweighted samples, that is, events that are distributed according to the probability distribution that describes the physical interactions. Second, we re-compute the amplitude values corresponding to these phase-space samples with MadGraph’s standalone module.
@@
For the L-GATr model, we embed each particle as a token. It is characterized with its four-momentum, embedded as a grade-1 multivector in the geometric algebra, as well as a one-hot embedding of the particle type, embedded as a scalar. We standardize the four-momentum inputs to unit variance, using the same normalization for each component to preserve Lorentz equivariance. In addition to the particle tokens we use one “global” token, initialized to zero. After processing these inputs with an L-GATr network, we select the scalar component of the global token and identify it as the amplitude output. We use 8 attention blocks, 32 multivector and 32 scalar channels, and 8 attention heads, resulting in $1.8\times 10^{6}$ learnable parameters.
For the CGENN, we minimally alter the graph neural network version of the model built for the top tagging classification task so that it is able to perform amplitude regression. We keep the hyperparameters proposed by [68] and use 72 hidden node features, 8 hidden edge features, and 4 blocks. This model features around $3.2\times 10^{5}$ trainable parameters.
@@
All models are trained by minimizing a mean squared error (MSE) loss on the preprocessed amplitude targets and by making use of the Adam optimizer. We use a batch size of 256 and a fixed learning rate of $10^{-4}$ for all baselines. As for the number of training steps, MLP and DSI are trained for around $2.5\times 10^{6}$ iterations, the Transformer for around $10^{6}$ iterations and GAP, CGENN and GATr for $2.5\times 10^{5}$ iterations. We use no regularization method for any of our baselines. We use early stopping across all training runs.
@@ C.2 Top tagging
As the top-tagging datasets operate with reconstructed particles as measured by a detector, we include the proton beam direction as an extra particle input to the network, which partially breaks the Lorentz equivariance of the process. We encode it as a rank-2 multivector representing the plane orthogonal to the beam direction. We also add another token containing the time direction $(1,0,0,0)$ as a rank-1 multivector. The time direction is required to break the special orthochronous Lorentz group $SO^{+}(1,3)$ down to the subgroup $SO(3)$.
@@
L-GATr is trained by minimizing a binary cross entropy (BCE) loss on the top quark labels. We train it for $2\times 10^{5}$ steps using the EvoLved Sign Momentum (LION) [27] optimizer with a weight decay of 0.2 and a batch size of 128. We use a Cosine Annealing scheduler [56] with the maximum learning rate set at $3\times 10^{-4}$.